# A Question-centric Multi-experts Contrastive Learning Framework for Improving the Accuracy and Interpretability of Deep Sequential Knowledge Tracing Models

### v1

@@ 2.2. Interpreting Deep Learning Based Knowledge Tracing Models
In recent years, an increasing number of interpretable techniques have been integrated into deep learning knowledge tracing (DLKT) models, enhancing both student modeling and prediction tasks. These techniques can be categorized into three distinct categories, as follows:
  • C1: Post-hoc local explanation. Post-hoc local explanation techniques are employed in the knowledge tracing (KT) task to scrutinize each individual prediction and understand the underlying reasoning behind the decisions made by DLKT models (Lu et al., 2020, 2022). For example, Lu et al. (2022) utilize a layer-wise relevance propagation method, which leverages the back propagating relevance scores from the model’s output layer to interpret the deep sequential KT model.
  • C2: Global interpretability with explainable structures. Aggregate an interpretable cognitive module into existing deep learning knowledge tracing (DLKT) architectures to provide a transparent and interpretable representation of the student’s knowledge state (Wang et al., 2020b; Zhao et al., 2020; Pu et al., 2022). For example, Wang et al. (2020b) incorporate an intermediate interaction layer based on multidimensional Item Response Theory (IRT) and model both student factors and exercise factors, allowing for a more fine-grained understanding of the knowledge state modeling process. Pu et al. (2022) introduce an automatic temporal cognitive method to enhance the modeling of changes in students’ knowledge states over time.
  • C3: Global interpretability with explainable parameters. Get the explainable parameters obtained from outputs of the DLKT models and directly feed it into cognitively interpretable models to estimate the probability that a student will answer a question correctly (Converse et al., 2021; Yeung, 2019b; Chen et al., 2023). For example, Chen et al. (2023) cast the parameters from problem solving and question-agnostic knowledge mastery modules into the IRT-like form. Yeung (2019b) model the levels of student abilities and KC difficulties with a dynamic key-value memory network for KT task and feed the learned results to an IRT layer for final prediction.
@@ 2.3. Constrastive Learning
Contrastive learning incorporates two major procedures: the construction of positive and negative pairs and the calculation of the contrastive loss. The contrastive loss function measures the similarity between the representations of positive pairs and encourages their embeddings to be close in the feature space. Simultaneously, it enforces the representations of negative pairs to be distant from each other. This process assists model in learning more discriminative and informative representations. As contrastive learning has achieved great success in the realm of representation learning, its application has extended to various downstream scenarios. For example, Gao et al. (2021) use a dropout trick to derive positive samples in the embedding level, and then apply both supervised and self-supervised methods to acquire better sentence embedding. Radford et al. (2021); Li et al. (2020) use contrastive learning to pre-train a vision language model to align the message between images and their corresponding text.  Zhang et al. (2022) regard word and definition as a semantic equivalence pair to do contrastive learning to enhance the word and definition representations. Zhang et al. (2023) implement contrastive learning as an auxiliary task to obtain a more oriented and pure task prompt representation. Gunel et al. (2020) design an auxiliary contrastive loss to enhance the generalization performance of rare examples and improve the robustness of the model. In the KT task, Lee et al. (2022) propose CL4KT framework. They do automatic data augmentation (e.g. question masking, interaction cropping, and interaction permutation) on the original student learning sequence to obtain the positive and negative pairs of learning sequence. Subsequently, it pools these sequence representation into an overall dense embedding to conduct contrastive learning. However, the positive and negative pairs are built at the learning sequence level rather than the direct question level, and the pairs generated by the automatic data augmentation method lack the necessary semantic alignment information, which can lead to inaccuracies and inconsistencies when doing contrastive learning. Furthermore, the pooling operation used to acquire representations for computing batch contrastive loss results in the loss of sequence information, thereby adversely affecting question-level representation learning during contrastive learning. In this work, we carefully design question-centric positive and negative pairs that encompass semantic alignment information at a fine-grained level. Additionally, our proposed method directly learns question-level representations without relying on pooling operations, making the learning procedure more effective. This thoughtful design aims to enhance the representation of the less interactive question, thereby leading to a improvement in accurately getting its question knowledge acquisition states for interpretable prediction layer.
@@ 2.4. Mixture of Experts
The mixture of experts technique comprises two key components: multiple expert networks and a gating network. The expert networks focus on capturing patterns or information from different perspectives, while the gating network acts as a decision-maker, determining the importance or contribution of each expert’s output. By combining the expertise of different expert networks, the mixture of experts technique can leverage the strengths of each expert to enhance the overall performance. Recently, the mixture of experts technique has gained significant popularity in various domains, such as computer vision, natural language processing, and rcommender systems (Lepikhin et al., 2020; Fedus et al., 2022; Lewis et al., 2021; Du et al., 2022; Artetxe et al., 2021; Ryabinin and Gusev, 2020; Riquelme et al., 2021; Liang et al., 2021; Alshaikh et al., 2020). Since the mixture-of-experts is very conducive to parallelization, some works have taken this technique to build the large-scale language and vision models (Lepikhin et al., 2020; Fedus et al., 2022; Lewis et al., 2021; Du et al., 2022; Artetxe et al., 2021; Ryabinin and Gusev, 2020; Riquelme et al., 2021). By employing a gating network to select the most suitable expert during the forward pass, these models can scale up their parameters without incurring additional computational overhead. Moreover, the mixture of experts has been leveraged to obtain more comprehensive and robust representations or prediction results (Liang et al., 2021; Alshaikh et al., 2020; Jacobs et al., 1991). In our work, we adopt the mixture of experts technique to enhance the robustness and accuracy of the knowledge acquisition states for the interpretable prediction layer.
@@ 4.1. Interaction Encoder
Interaction projection methods such as qDKT (Sonkar et al., 2020), which neglect the relationships between questions and concepts (KCs) and rely solely on the question sequence to track students’ knowledge states, may not provide an adequate representation of interaction.
Therefore, we follow Long et al. (2021b) to encode the interaction in a more granular manner by capturing the relationship between questions and concepts (KCs). In our methodology, we employ separate raw interaction encodings $\mathbf{e}_{t}$ and $\mathbf{c}_{t}$ for the MQKA module and the MCKA module respectively. Specifically, in the MQKA module, we represent each question level interaction $\mathbf{e}_{t}$ as a concatenation of the question, the student’s response, and the corresponding set of KCs associated with the current question, i.e.,
@@ 4.3. Multi-experts Concept Knowledge Acquisition Module
However, in real educational settings, students often make guesses or slip-ups when interacting with certain questions. This can lead to the MQKA module becoming overly sensitive to individual interactions, resulting in inaccuracies during the inference stage. To mitigate the issue of inaccuracies caused by guesses or slip-ups, we introduce a supplemental knowledge acquisition, namely concept knowledge acquisition. Concept knowledge acquisition can be regarded as a more general acquisition of knowledge, because the interactions of student in specific concept is much more than in specific question3.
To model concept knowledge acquisition, we propose the MCKA (Multi-experts Concept Knowledge Acquisition) module. Similar to the question knowledge acquisition modeling in the MQKA module, we utilize an additional LSTM cell to update the student’s concept knowledge state $\mathbf{v}_{t}$ after receiving each response. The LSTM cell in the MCKA module takes concept-level interaction $\mathbf{c}_{t}$ instead of $\mathbf{e}_{t}$:
@@ 4.6. Optimization of Q-MCKT
$\mathcal{L}_{\mbox{\tiny KT}}$ and *$\mathcal{L}_{*}(\mathbf{z})$ are defined as follows:
@@
*$\mathcal{L}_{*}(\mathbf{z})=-\sum_{i}\bigl{(}r_{i}\log\sigma(z_{i})+(1-r_{i})\log(1-\sigma(z_{i}))\bigl{)}$
@@ 5. Experiments
In this section, we present details of our experiment settings and the corresponding results. We conduct comprehensive analysis and investigations to illustrate the effectiveness of our Q-MCKT model. Specifically, we aim to answer the following research questions: (RQ1) How does the proposed Q-MCKT framework perform compared to the state-of-the-art KT methods? (RQ2) What is the influence of components in Q-MCKT, such as the interpretable prediction layer and concept knowledge acquisition module? (RQ3) How does the CL affect the performance of Q-MCKT? (RQ4) How does the number of experts influence the performance of Q-MCKT? (RQ5) Can Q-MCKT model accurately estimate the knowledge acquisition state?
@@ 5.1. Datasets
  • ASSISTments2009121212https://sites.google.com/site/assistmentsdata/home/2009-2010-assistment-data/skill-builder-data-2009-2010 (ASSIST2009): is collected from ASSISTment online tutoring platform in the school year 2012-2013 that students are assigned to answer similar exercises from the skill builder problem sets.
  • Algebra 2005-2006131313https://pslcdatashop.web.cmu.edu/KDDCup/ (Algebra2005): stems from KDD Cup 2010 EDM Challenge which includes 13-14 year-old students’ interactions with Algebra questions. It has detailed step-level student responses to the mathematical problems.
@@ 5.2. Baselines
  • DKT+: an improved version of DKT to solve the reconstruction and non-consistent prediction problems (Yeung and Yeung, 2018b).
  • DKT-F: an extension of DKT that model the students’ forgetting behavior on the DKT model (Nagatani et al., 2019a).
@@
  • SparseKT: employs a soft-thresholding sparse attention technique to mitigate the risk of overfitting (Huang et al., 2023).
@@ 5.3. Experimental Setup
To evaluate the model’s performance, we perform 5-fold cross-validation for every combination of models and datasets. We set the maximum length of model input sequence to 200. We use 80% of student sequences for training and validation, and use the rest 20% of student sequences for model evaluation. We adopt Adam optimizer (Kingma and Ba, 2015) to train all the models. The number of training epochs is set to 200. We choose to use an early stopping strategy that stops optimization when the AUC score is failed to get the improvement on the validation set in the latest 10 epochs. We adopt the Bayesian search method to find the best hyper-parameters for each fold by using wandb tool [^16] https://wandb.ai/. Specifically, the hyper-parameter $\lambda_{1}$ and $\lambda_{2}$, the learning rate and the embedding size $d$ are searching from [0,0.5,1,1.5,2], [1e-3, 1e-4, 1e-5], [64, 256] respectively. Following all existing DLKT research (Liu et al., 2023b; Guo et al., 2021; Ghosh et al., 2020; Choi et al., 2020a), we use the Area Under the Curve (AUC) as the main evaluation metric. We also choose to use Accuracy as the secondary evaluation metric.
@@ 5.4. Overall Performance (RQ1)
Table 2. The AUC and Accuracy performance of all the baseline models and our proposed model. We highlight the highest results with bold. Marker ***, $\circ$, and $\bullet$ indicates whether our proposed model is statistically superior/equal/inferior to the compared method (using paired t-test at 0.01 significance level).
@@
SimpleKT 0.7744±0.0018* 0.8254±0.0003* 0.8035±0.0000$\circ$ 0.6599±0.0027* 0.7320±0.0012* 0.8083±0.0005* 0.7328±0.0001* 0.6557±0.0029*
SparseKT 0.7739±0.0005* 0.8152±0.0030* 0.8033±0.0008* 0.6628±0.0036* 0.7282±0.0016$\bullet$ 0.8017±0.0020* 0.7322±0.0012* 0.6556±0.0048*
Ours 0.7876±0.0016 0.8429±0.0011 0.8035±0.0004 0.7228±0.0023 0.7376±0.0011 0.8239±0.0006 0.7329±0.0005 0.7067±0.0008
@@
The overall model performance is reported in Table 2. From Table 2, we make the following observations: (1) Our proposed model significantly outperforms 15 baselines on all four datasets in terms of AUC metric, and only has one loss with AKT in ASSIST2009 dataset in terms of Accuracy metric. More importantly, as a representative of the deep sequential KT models, compared with DKT, our proposed model improves the AUC by 3.30%, 2.80%, 3.46%, and 10.95% on four datasets. That shows our proposed method can significantly improve the performance. (2) When comparing performance on ASSIST2009, Algebra2005 to NIPS34 and EdNet, DLKT models behave quite differently. For example, DKT significantly outperforms qDKT on the ASSIST2009 and Algebra2005 datasets in AUC by 5.30% and 6.60% but is beaten by qDKT by 3.10% and 8.54% on the NIPS34 and EdNet datasets. Meanwhile, SAINT performs terrible in ASSIST2009 and Algebra2005 datasets, but is good on NIPS34 and EdNet datasets. We argue that it is because the average number of interactions per question (Avg .Q) for the ASSIST2009 and Algebra2005 are smaller than that for the NIPS34 and EdNet datasets (As we can see from Table 1, the values of Avg. Q are 12.9 and 3.8 in ASSIST2009 and Algebra2005 datasets). This results in the question-centric modeling overfitting to the data, which adversely affects predictive accuracy. (3) Results between DeepIRT and DKVMN are very close on three datasets, which empirically shows that the IRT function won’t sacrifice the model prediction ability too much. (4) AKT is a very strong baseline, it outperforms almost all other baseline methods on all datasets. Since it learns the Rasch model-based embeddings that implicitly model question difficulties, it further empirically verifies the importance of considering question representations when building the DLKT models. (5) Our proposed Q-MCKT exhibits a more substantial improvement over the baselines on the EdNet dataset compared to the other three datasets. We attribute this to the fact that the EdNet dataset has a larger number of average KCs per question (Avg. per Q = 2.245), indicating that the majority of questions in EdNet encompass more than two KCs. However, most baseline methods encode the student’s learning sequence at the concept level, utilizing concepts to index questions. In this process, each question will be transformed into several KC interactions, and questions covering the same KC are treated as identical questions. This approach disregards much valuable question-centric information and introduces noise in the encoding process. This is particularly pronounced in datasets with larger Avg. per Q, such as EdNet, where questions containing more KCs result in a transformed process that includes additional noise. In contrast, as illustrated in Section 4.1, our Q-MCKT adopts a different approach by encoding the student’s learning sequence at the question level. This allows for the capture of both question- and concept-centric information in a more granular manner. Consequently, this approach effectively mitigates the aforementioned problems associated with datasets featuring larger Avg. per Q, as exemplified by EdNet.

### v2

@@ 2.2. Interpreting Deep Learning Based Knowledge Tracing Models
In recent years, an increasing number of interpretable techniques have been integrated into DLKT models, enhancing both student modeling and prediction tasks. These techniques can be categorized into three distinct categories, as follows:
  • C1: Post-hoc local explanation. Post-hoc local explanation techniques are employed in the KT task to scrutinize each individual prediction and understand the underlying reasoning behind the decisions made by DLKT models (Lu et al., 2020, 2022). For example, Lu et al. (2022) utilized a layer-wise relevance propagation method, which leveraged the back propagating relevance scores from the model’s output layer to interpret the deep sequential KT model.
  • C2: Global interpretability with explainable structures. Aggregate an interpretable cognitive module into existing DLKT architectures to provide a transparent and interpretable representation of the student’s knowledge state (Wang et al., 2020b; Zhao et al., 2020; Pu et al., 2022). For example, Wang et al. (2020b) incorporated an intermediate interaction layer based on multidimensional IRT and model both student factors and exercise factors, allowing for a more fine-grained understanding of the knowledge state modeling process. Pu et al. (2022) introduced an automatic temporal cognitive method to enhance the modeling of changes in students’ knowledge states over time.
  • C3: Global interpretability with explainable parameters. Get the explainable parameters obtained from outputs of the DLKT models and directly feed it into cognitively interpretable models to estimate the probability that a student will answer a question correctly (Converse et al., 2021; Yeung, 2019b; Chen et al., 2023). For example, Chen et al. (2023) cast the parameters from problem solving and question-agnostic knowledge mastery modules into the IRT-like form. Yeung (2019b) modeled the levels of student abilities and KC difficulties with a dynamic key-value memory network for KT task and fed the learned results to an IRT layer for final prediction.
@@ 2.3. Constrastive Learning
Contrastive learning incorporates two major procedures: the construction of positive and negative pairs and the calculation of the contrastive loss. The contrastive loss function measures the similarity between the representations of positive pairs and encourages their embeddings to be close in the feature space. Simultaneously, it enforces the representations of negative pairs to be distant from each other. This process assists model in learning more discriminative and informative representations. As contrastive learning has achieved great success in the realm of representation learning, its application has extended to various downstream scenarios. For example, Gao et al. (2021) used a dropout trick to derive positive samples in the embedding level, and then applied both supervised and self-supervised methods to acquire better sentence embedding. Radford et al. (2021); Li et al. (2020) used contrastive learning to pre-train a vision language model to align the message between images and their corresponding text.  Zhang et al. (2022) regarded word and definition as a semantic equivalence pair to do contrastive learning to enhance the word and definition representations. Zhang et al. (2023) implemented contrastive learning as an auxiliary task to obtain a more oriented and pure task prompt representation. Gunel et al. (2020) designed an auxiliary contrastive loss to enhance the generalization performance of rare examples and improve the robustness of the model. In the KT task, Lee et al. (2022) proposed CL4KT framework. They did automatic data augmentation (e.g. question masking, interaction cropping, and interaction permutation) on the original student learning sequence to obtain the positive and negative pairs of learning sequence. Subsequently, it pooled these sequence representation into an overall dense embedding to conduct contrastive learning. However, the positive and negative pairs are built at the learning sequence level rather than the direct question level, and the pairs generated by the automatic data augmentation method lack the necessary semantic alignment information, which can lead to inaccuracies and inconsistencies when doing contrastive learning. Furthermore, the pooling operation used to acquire representations for computing batch contrastive loss results in the loss of sequence information, thereby adversely affecting question-level representation learning during contrastive learning. In this work, we carefully design question-centric positive and negative pairs that encompass semantic alignment information at a fine-grained level. Additionally, our proposed method directly learns question-level representations without relying on pooling operations, making the learning procedure more effective. This thoughtful design aims to enhance the representation of the less interactive question, thereby leading to a improvement in accurately getting its question knowledge acquisition states for interpretable prediction layer.
@@ 2.4. Mixture of Experts
The mixture of experts technique comprises two key components: multiple expert networks and a gating network. The expert networks focus on capturing patterns or information from different perspectives, while the gating network acts as a decision-maker, determining the importance or contribution of each expert’s output. By combining the expertise of different expert networks, the mixture of experts technique can leverage the strengths of each expert to enhance the overall performance. Recently, the mixture of experts technique has gained significant popularity in various domains, such as computer vision, natural language processing, and recommender systems (Lepikhin et al., 2020; Fedus et al., 2022; Lewis et al., 2021; Du et al., 2022; Artetxe et al., 2022; Ryabinin and Gusev, 2020; Riquelme et al., 2021; Liang et al., 2021; Alshaikh et al., 2020). Since the mixture-of-experts is very conducive to parallelization, some works have taken this technique to build the large-scale language and vision models (Lepikhin et al., 2020; Fedus et al., 2022; Lewis et al., 2021; Du et al., 2022; Artetxe et al., 2022; Ryabinin and Gusev, 2020; Riquelme et al., 2021). By employing a gating network to select the most suitable expert during the forward pass, these models can scale up their parameters without incurring additional computational overhead. Moreover, the mixture of experts has been leveraged to obtain more comprehensive and robust representations or prediction results (Liang et al., 2021; Alshaikh et al., 2020; Jacobs et al., 1991). In our work, we adopt the mixture of experts technique to enhance the robustness and accuracy of the knowledge acquisition states for the interpretable prediction layer.
@@ 4.1. Interaction Encoder
Interaction projection methods such as qDKT (Sonkar et al., 2020), which neglect the relationships between questions and KCs and rely solely on the question sequence to track students’ knowledge states, may not provide an adequate representation of interaction.
Therefore, we follow Long et al. (2021) to encode the interaction in a more granular manner by capturing the relationship between questions and KCs. In our methodology, we employ separate raw interaction encodings $\mathbf{e}_{t}$ and $\mathbf{c}_{t}$ for the MQKA module and the MCKA module respectively. Specifically, in the MQKA module, we represent each question level interaction $\mathbf{e}_{t}$ as a concatenation of the question, the student’s response, and the corresponding set of KCs associated with the current question, i.e.,
@@ 4.3. Multi-experts Concept Knowledge Acquisition Module
However, in real educational settings, students often make guesses or slip-ups when interacting with certain questions. This can lead to the MQKA module becoming overly sensitive to individual interactions, resulting in inaccuracies during the inference stage. To mitigate the issue of inaccuracies caused by guesses or slip-ups, we introduce a supplemental knowledge acquisition, namely concept knowledge acquisition. Concept knowledge acquisition can be regarded as a more general acquisition of knowledge, because the interactions of student in specific concept is much more than in specific question.
To model concept knowledge acquisition, we propose the multi-experts concept knowledge acquisition module. Similar to the question knowledge acquisition modeling in the MQKA module, we utilize an additional LSTM cell to update the student’s concept knowledge state $\mathbf{v}_{t}$ after receiving each response. The LSTM cell in the MCKA module takes concept-level interaction $\mathbf{c}_{t}$ instead of $\mathbf{e}_{t}$:
@@ 4.6. Optimization of Q-MCKT
$\mathcal{L}_{\mbox{\tiny KT}}$ and $\mathcal{L}_{*}(\mathbf{z})$ are defined as follows:
@@
$\mathcal{L}_{*}(\mathbf{z})=-\sum_{i}\bigl{(}r_{i}\log\sigma(z_{i})+(1-r_{i})\log(1-\sigma(z_{i}))\bigl{)}$
@@ 5. Experiments
In this section, we present details of our experiment settings and the corresponding results. We conduct comprehensive analysis and investigations to illustrate the effectiveness of our Q-MCKT model. Specifically, we aim to answer the following research questions: (RQ1) How does the proposed Q-MCKT framework perform compared to the state-of-the-art KT methods? (RQ2) What is the influence of components in Q-MCKT, such as the interpretable prediction layer and concept knowledge acquisition module? (RQ3) How does the CL affect the performance of Q-MCKT? (RQ4) How does the number of experts influence the performance of Q-MCKT? (RQ5) Can Q-MCKT model accurately estimate the knowledge acquisition state? (RQ6) How does Q-MCKT produce interpretable prediction results?
@@ 5.1. Datasets
  • ASSISTments2009777https://sites.google.com/site/assistmentsdata/home/2009-2010-assistment-data/skill-builder-data-2009-2010 (ASSIST2009): is collected from ASSISTment online tutoring platform in the school year 2012-2013 that students are assigned to answer similar exercises from the skill builder problem sets.
  • Algebra 2005-2006888https://pslcdatashop.web.cmu.edu/KDDCup/ (Algebra2005): stems from KDD Cup 2010 EDM Challenge which includes 13-14 year-old students’ interactions with Algebra questions. It has detailed step-level student responses to the mathematical problems.
@@ 5.2. Baselines
  • DKT+: an improved version of DKT to solve the reconstruction and non-consistent prediction problems (Yeung and Yeung, 2018).
  • DKT-F: an extension of DKT that model the students’ forgetting behavior on the DKT model (Nagatani et al., 2019).
@@ 5.3. Experimental Setup
To evaluate the model’s performance, we perform 5-fold cross-validation for every combination of models and datasets. We set the maximum length of model input sequence to 200. We use 80% of student sequences for training and validation, and use the rest 20% of student sequences for model evaluation. We adopt Adam optimizer (Kingma and Ba, 2015) to train all the models. The number of training epochs is set to 200. We choose to use an early stopping strategy that stops optimization when the AUC score is failed to get the improvement on the validation set in the latest 10 epochs. We adopt the Bayesian search method to find the best hyper-parameters for each fold by using wandb tool [^11] https://wandb.ai/. Specifically, the hyper-parameter $\lambda_{1}$ and $\lambda_{2}$, the learning rate and the embedding size $d$ are searching from {0, 0.5, 1, 1.5, 2}, {1e-3, 1e-4, 1e-5}, {64, 256} respectively. Following all existing DLKT research (Liu et al., 2023b; Guo et al., 2021; Ghosh et al., 2020; Choi et al., 2020a), we use the Area Under the Curve (AUC) as the main evaluation metric. We also choose to use Accuracy as the secondary evaluation metric.
@@ 5.4. Overall Performance (RQ1)
Table 2. The AUC and Accuracy performance of all the baseline models and our proposed Q-MCKT model. We highlight the highest results with bold. Marker *, $\circ$, and $\bullet$ indicates whether our proposed model is statistically superior/equal/inferior to the compared baseline models (using paired t-test at 0.01 significance level).
@@
SimpleKT 0.7744±0.0018* 0.8254±0.0003* 0.8035±0.0000$\circ$ 0.6599±0.0027* 0.7320±0.0012* 0.8083±0.0005* 0.7328±0.0001$\circ$ 0.6557±0.0029*
Q-MCKT 0.7876±0.0016 0.8429±0.0011 0.8035±0.0004 0.7228±0.0023 0.7376±0.0011 0.8239±0.0006 0.7329±0.0005 0.7067±0.0008
@@
The overall model performance is reported in Table 2. From Table 2, we make the following observations: (1) Our proposed model significantly outperforms 14 baselines on all four datasets in terms of AUC metric, and only has one loss with AKT in ASSIST2009 dataset in terms of Accuracy metric. More importantly, as a representative of the deep sequential KT models, compared with DKT, our proposed model improves the AUC by 3.30%, 2.80%, 3.46%, and 10.95% on four datasets. That shows our proposed method can significantly improve the performance. (2) When comparing performance on ASSIST2009, Algebra2005 to NIPS34 and EdNet, DLKT models behave quite differently. For example, DKT significantly outperforms qDKT on the ASSIST2009 and Algebra2005 datasets in AUC by 5.30% and 6.60% but is beaten by qDKT by 3.10% and 8.54% on the NIPS34 and EdNet datasets. Meanwhile, SAINT performs terrible in ASSIST2009 and Algebra2005 datasets, but is good on NIPS34 and EdNet datasets. We argue that it is because the average number of interactions per question (Avg .Q) for the ASSIST2009 and Algebra2005 are smaller than that for the NIPS34 and EdNet datasets (As we can see from Table 1, the values of Avg. Q are 12.9 and 3.8 in ASSIST2009 and Algebra2005 datasets). This results in the question-centric modeling overfitting to the data, which adversely affects predictive accuracy. (3) Results between DeepIRT and DKVMN are very close on three datasets, which empirically shows that the IRT function won’t sacrifice the model prediction ability too much. (4) AKT is a very strong baseline, it outperforms almost all other baseline methods on all datasets. Since it learns the Rasch model-based embeddings that implicitly model question difficulties, it further empirically verifies the importance of considering question representations when building the DLKT models. (5) Our proposed Q-MCKT exhibits a more substantial improvement over the baselines on the EdNet dataset compared to the other three datasets. We attribute this to the fact that the EdNet dataset has a larger number of average KCs per question (Avg. per Q = 2.245), indicating that the majority of questions in EdNet encompass more than two KCs. However, most baseline methods encode the student’s learning sequence at the concept level, utilizing concepts to index questions. In this process, each question will be transformed into several KC interactions, and questions covering the same KC are treated as identical questions. This approach disregards much valuable question-centric information and introduces noise in the encoding process. This is particularly pronounced in datasets with larger Avg. per Q, such as EdNet, where questions containing more KCs result in a transformed process that includes additional noise. In contrast, as illustrated in Section 4.1, our Q-MCKT adopts a different approach by encoding the student’s learning sequence at the question level. This allows for the capture of both question- and concept-centric information in a more granular manner. Consequently, this approach effectively mitigates the aforementioned problems associated with datasets featuring larger Avg. per Q, as exemplified by EdNet.